• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Hall A (Techs)

    Log entry time 15:55:14 on December 09, 2014

    Entry number 404726


    "Checklist complete and target window and spectrometer guards are removed"

    "Angle limits for the left spectrometer are 36.5 degrees to 36.5 degrees"

    "Angle limits for the right spectrometer are 67.25 degrees to 67.25 degrees"

    "The tech on call is HEIDI FANSLER"

    "Sieve Slit is "NOT" installed on right/left spectrometers"

    "Right spectrometer window and VDC covers are "NOT" removed, see Jack for more information"

    "See jack for powering up magnets on right spectrometer"

    Lead flows are at diminished capacity per Heidi will update status later"

    "Left dipole still being worked on by Jack Segal"